# Embodying Physical Computing into Soft Robots

### v1

@@ Introduction
The dream of creating entirely soft, versatile, and capable robots — akin to the octopus — has long inspired scientists and engineers. We have witnessed significant progress in soft actuation (?, ?), sensing (?), and power (?), enabling these robots to operate in a wide range of challenging environments, from deep within our own bodies (?) to the far bottom of the Mariana trench (?). Yet, softening and onboarding computers and controllers remain a major challenge and present one of the final frontiers towards robust and intelligent soft robots suitable for everyday use. In this regard, roboticists have long recognized that the inherent material softness can facilitate and simplify control, and many innovative strategies have been explored. For example, soft and rotating legs can naturally accommodate uneven surfaces and large obstacles, allowing the robot to traverse challenging terrains without complex controls like in the quadrupeds  (?). Soft curling tentacles can wrap and entangle themselves around objects with widely different shapes, thus manipulating them with a simple global pressure input  (?). Such softness-facilitated control is sometimes referred to as “intelligent my mechanics” (?) or “morphological computation” (?, ?). They offer exciting potential, but frequently lack the sophistication and (re-) programmability available from the more universal controllers based on digital computation.
@@
In parallel with the advancements in soft robotics, there is also growing interest in CMOS-free physical computers  (?, ?, ?, ?, ?). This emerging paradigm seeks to encode inputs into a mechanical construct physically — for example, in the form of deformation   (?, ?, ?, ?), fluid flow (?, ?), thermal heat flux (?), or waves  (?, ?, ?, ?, ?) — and leverage the internal interactions among its constituent elements to process these inputs according to a programmed evolution. The resulting output typically remains in the same physical domain as the input so that it can be easily decoded and interpreted. This idea of performing computation without the CMOS electronics could benefit us with a higher energy efficiency, massive parallelization, and resiliency against adversarial working conditions. To realize physical computing, researchers have explored two main approaches. One borrows the analog computing concept and uses dynamic and continuously variable signals to perform computation (e.g., using acoustic waves to solve differential equations (?, ?) or using mechanical vibrations to perform machine learning tasks (?, ?)). The other strategy borrows the algorithmic computing concepts and uses discrete mechanisms to construct logic gates and complex physical circuitry (e.g., using mechanical AND, OR, and NOR logic gates to construct a full adder (?)).
@@
Therefore, there is a tremendous opportunity to converge the two fields of soft robotics and physical computing. That is, one could construct a physical computer out of soft materials and integrate it with soft sensors and actuators. Such integration can lead to a new class of entirely soft computation and control methods with flexibility and programmability for more sophisticated robotic tasks. As a result, a new discipline can emerge to bridge material science, information theory, and soft robotic control. Over the past several years, we have witnessed a rapid emergence of soft robots with integrated and embodied physical computers. Therefore, this perspective paper aims to capture this momentum and establish a more formal framework for embodying physical computing in soft robots. Consequently, it will survey the different analog and algorithmic computers implemented in the soft robotic body and discuss future directions.
@@ 1 What is physical computing (and what is not)?
Therefore, in this paper, computing does not exist without encoding, decoding, and programming (?). Under this formalism, some nonconventional and innovative computing paradigms in the robotics field, such as the aforementioned “morphological computation,” are beyond our scope. On the other hand, a mechanical construct — e.g., architected materials or soft robotic body — that can incorporate encoding, decoding, and programming would meet the necessary condition to function as a physical computer.
Moreover, based on these definitions, we will adopt the theory from Jaeger et. al, and categorize physical computers into two sets (Figure 1b) (?, ?).
@@
Figure 1: The architecture and categorization of physical computing. (a) The computing architecture adopted in this perspective includes input encoding, output decoding, and programmable input-output evolution. (b-d) The two types of physical computing are analog (up) and algorithmic (bottom). (b) The early attempts at analog and algorithmic computing were purely mechanical (the Harmonic Analyzer photo adapted from (?) with permission, and the Difference Machine photo credits to Science Museum London, $C$ Attribution-SA 2.0). (c) The electronic implementations of these two computing approaches have become pillars of our modern life. (d) Examples of physical computing embodied in soft robots, including, from top to bottom, an electronics-free legged robot with an analog oscillator (?) (image credit to David Baillot, Jacobs School of Engineering, UC San Diego), a modular manipulator with embodied reservoir computing (image adapted from (?) $C$ BY 4.0), and a robotic hand with fluidic logic control (image adaptive from (?) with permission).
@@
One is analog, where input and output signals are continuous, and the evolution from input to output is governed by smooth (and frequently physics-based) functions. Albert Michelson’s harmonic analyzer (?) and our human neural system are classical examples of analog computers. In soft robots, this can be accomplished by exploiting their bodies’ nonlinear dynamic responses for physical reservoir computing. The other type of physical computer is the algorithmic, where the input and output take a discrete format, and the evolution from input to output is programmed via a set of abstract logical rules. Charles Babbage’s difference machine (?) and our omnipresent CMOS-based computing chips are classical examples of an algorithmic computer. In soft robots, this can be accomplished by, for example, an assembly of mechanical Boolean gates featuring elastic bistability (i.e., mechanical logic gates).
@@ 2 Analog Oscillators and Rhythmic Motion
Figure 2: Analog oscillators and rhythmic deformation. (a) Electronics-free pneumatic control: A soft ring-oscillator circuit generates rhythmic leg actuation from a constant pressure input, enabling a quadruped to produce diagonal-couplet walking gaits. A bistable 4/2 switch selects gait direction, and dual oscillators set the phase between leg pairs (image adapted from  (?) with permission). (b) Controller-free SMA modular robot: A curved mono stable beam and a mechanical slider can convert a single DC power supply into sustained self-oscillation. And a bistable switch can alternate power supply between the front and back modules for out-of-phase deformation and crawling (image adapted from  (?) with permission). (c) Twisted LCE ribbon robot: The ambient heating drives continuous self-rolling of this robotic structure for locomotion. When the robot contacts an obstacle, it will store elastic energy and then snaps to reverse its direction, enabling autonomous avoidance and maze escape (image adapted from  (?) with permission).
For example, in the quadruped robot shown in Figure 2(a), rhythmic and coordinated leg swing motions are generated by an entirely pneumatic ring oscillator  (?). In this oscillator, a small pressure tank supplies a constant pressure ($P_{+}$) to drive the analog oscillator (input encoding). Inside the oscillator circuit, three soft valves serve as inverters with a built-in delay, and a snap-through membrane alternates between closed and open flow paths, allowing the high-pressure flow to advance around the ring. This mechanism essentially transforms the steady input into a phase-shifted sequence of pressure pulses at the three nodes (programmed input-output evolution). Finally, the pulsed pneumatic pressure flows to the corresponding soft legs, which convert the pressure inflation into mechanical swing motions for walking (output decoding). In addition, a soft bistable valve and tethered mechanical controller are added to swap two connections like a latching switch, so triggering the valve can reverse the output pressure pulse sequence, thus reversing the locomotion direction (re-programming).
@@ 3 Analog Physical Reservoir Computing
While an analog oscillator provides a promising alternative to micro-controllers for computing and generating rhythmic motions, its information processing capability is hard-coded and fixed by its structural design. In contrast, there is an emerging notion called informational embodiment, where the combinatorial richness of a soft robot’s body deformation can encode spatiotemporal patterns without relying on symbolic or centralized representations   (?). This concept bridges soft body dynamics and computation, offering another pathway toward decentralized and analog computing.
@@
Figure 3: Open and closed-loop physical reservoir computing in soft robots. (a) Open-loop PRC for information perception: A modular manipulator with embedded strain gauges is driven by SMA actuators. Its high-dimensional body dynamics (measured by the strain $s_{i}(t)$) serve as reservoir states, which can be processed with trained linear readout $w_{i}$ to decode and identify the payload (image adapted from (?) with permission). (b) Other examples of information perception with open-loop PRC, including terrain classification (image adapted from (?) with permission), wind detection on a compliant membrane wing (image adapted from (?) with permission), and a self-sensing shape memory alloy actuator that could predict its end effector trajectory (image adapted from (?) with permission). (c) Closed-loop PRC for embodied control: A quadruped robot uses its compliant spine as the reservoir. The four outputs of the reservoir kernel are fed back to the leg actuators to generate robust and adaptable locomotion gaits (image adapted from (?) with permission). (d) Other examples of control embodiments with closed-loop PRC, including manipulation with a multi-segment continuum arm (image adapted from (?) with permission) and a surface-swimming robot (adapted from (?) with permission).
@@
This example clearly illustrates how the open-loop reservoir computing framework enables soft robots to conduct spatiotemporal filtering through their intrinsic deformation dynamics, allowing them to extract complex information without requiring dense sensor arrays or extensive digital processing. As a result, numerous studies have emerged utilizing different soft robotic platforms. For instance, a fabric-based soft manipulator can estimate joint bending angle and payload weight simultaneously using only a few distributed pressure sensors  (?). Contact dynamics in a soft arm allow for tactile sensing and object property estimation without electronic skins  (?). For environment monitoring, a brush-like flexible sensor encodes surface textures through passive contact  (?)(Fig.3b, up), while in aerial applications, a flapping-wing robot detects wind direction directly from wing deformation, eliminating the need for airflow sensors  (?)(Fig.3b, middle). A shape memory alloy actuator reservoir  (?) demonstrates the capability of predicting the future trajectory of its end effector under various driving signals (Fig. 3b, bottom). Collectively, these demonstrations show that soft robots with sparse, low-dimensional sensors can nonetheless achieve high-dimensional perception by exploiting their own body as the physical reservoir—making PRC a minimalist yet powerful strategy for embodied sensing.
@@
Beyond information perception, PRC enables soft robots to autonomously generate periodic and robust motions by embedding control into their intrinsic body dynamics. That is, instead of connecting robotic actuators to external digital controllers, one can feed the body reservoir’s output back to these actuators for real-time motor behavior control. In this case, the deformation dynamics of the robotic body and its interaction with the environment play a critical role. A representative example is the quadruped robot with its flexible spine serving as the reservoir computing kernel (?)(Fig. 3c). In this case, the linear readout layer performs weighted linear summations, mapping the internal force and strain of the flexible spine into four control commands, each for a motored leg. The readouts are first trained in an open-loop setup with teacher forcing. Once the training is complete, the loop is closed, and the robot can perform trotting, bounding, and turning with a strong ability to recover from disturbance. This embodied controller demonstrates how PRC converts compliant mechanics into an energy-efficient control system: the body remains unchanged, only the readout is “programmed,” and feedback routes it from perception to action.
@@
4 Algorithmic Physical Computing

### v2

@@ Introduction
The dream of creating entirely soft, versatile, and capable robots — akin to the octopus — has long inspired scientists and engineers. We have witnessed significant progress in soft actuation (?, ?), sensing (?), and power (?), enabling these robots to operate in a wide range of challenging environments, from deep within our own bodies (?) to the far bottom of the Mariana trench (?). Yet, softening and onboarding computers and controllers remain a major challenge and present one of the final frontiers towards robust and intelligent soft robots suitable for everyday use. In this regard, roboticists have long recognized that the inherent material softness can facilitate and simplify control, and many innovative strategies have been explored. For example, soft and rotating legs can naturally accommodate uneven surfaces and large obstacles, allowing the robot to traverse challenging terrains without complex controls like in the quadrupeds  (?). Soft curling tentacles can wrap and entangle themselves around objects with widely different shapes, thus manipulating them with a simple global pressure input  (?). Such softness-facilitated control is sometimes referred to as “intelligence by mechanics” (?) or “morphological computation” (?, ?). They offer exciting potential, but frequently lack the sophistication and (re-) programmability available from the more universal controllers based on digital computation.
@@
In parallel with the advancements in soft robotics (and partly inspired by the need for soft robotic computing and control), there is also growing interest in CMOS-free physical computers  (?, ?, ?, ?, ?). This emerging paradigm seeks to encode physical inputs into a mechanical construct (or kernel) — for example, in the form of deformation   (?, ?, ?, ?, ?), fluid flow (?, ?), thermal heat flux (?), or waves  (?, ?, ?, ?, ?) — and leverage the internal interactions among the kernel’s constituent elements to process these inputs according to a programmed evolution. The resulting output typically remains in the same physical domain as the input so that it can be easily decoded and interpreted. The paths to physical computing are diverse: One can use acoustic waves to solve differential equations (?, ?), re-purpose mechanical vibrations like neuron signals to perform machine learning tasks (?, ?, ?), or construct mechanical logic gates and physical circuitry to perform algorithmic operations (?). Overall, the idea of performing computation without CMOS electronics could benefit us with higher energy efficiency (?), parallelization (?), and resiliency against adversarial working conditions.
@@
Therefore, there is a tremendous opportunity to introduce physical computing into the field of soft robotics. That is, one could construct a physical computer out of soft materials and integrate it with soft sensors and actuators. Such integration can lead to a new class of entirely soft computation and control methods with flexibility, robustness, and programmability for more sophisticated tasks. As a result, we have witnessed a rapid emergence of soft robots with integrated and embodied physical computers over the past several years. And these physically computing robots have become an important part of recent reviews that offer a birds-eye overview on embodied intelligence (or mechanical intelligence, physical control) in robotics (?, ?, ?, ?).
On the other hand, we believe a separate, deeper dive into physical computing in soft robots can benefit the research community. Specifically, we aim to define physical computing using a rigorous framework, including encoding, decoding, and a (re-)programmable computing kernel, and build upon this definition to categorize physical computing into two distinct types: analog and algorithmic (more in the following Section 1). In this way, we can dissect and re-examine soft robotics through a different lens. We will also have a more systematic framework to introduce new physical computing concepts from other disciplines to robotics.
Therefore, this perspective paper will first establish a more formal framework for physical computing, then survey the different analog and algorithmic computers implemented in the soft robots, and discuss challenges and future directions in the end.
@@ 1 What is physical computing (and what is not)?
Figure 1: The architecture and categorization of physical computing. (a) The computing architecture adapted in this perspective includes input encoding, output decoding, and programmable input-output evolution. (b) Analog computing: The Harmonic Analyzer is an elegant example of analog computing from the 1800s (photo credit to the Nimitz Library, United States Naval Academy). In the modern electronic computing paradigm, Artificial Neural Networks also operate on analog (continuous) data. We will discuss two types of physical analog computing in soft robots: oscillator and reservoir. For example, an electronics-free legged robot uses an analog oscillator to walk (?) (photo credit to David Baillot, Jacobs School of Engineering, UC San Diego), and a modular manipulator uses embodied reservoir computing to classify payloads (image adapted from (?) $C$ BY 4.0). (c) Algorithmic computing: The Difference Machine is one of the first algorithmic computers (photo credits to Science Museum London, $C$ Attribution-SA 2.0). Modern CMOS-based computers are built exclusively on algorithmic Boolean logic. Here, we survey how mechanical logic is implemented in soft robots. For example, a robotic hand operates with fluidic logic control (image adapted from (?) $C$ BY 4.0).
Therefore, in this paper, computing does not exist without encoding, decoding, and programming (?). Under this formalism, some nonconventional and innovative computing paradigms in the robotics field, such as the aforementioned “morphological computation,” are beyond our scope. Morphological computation generally refers to the idea that a robot body’s shape, deformation, and dynamics can perform part of the “computation” needed for control. Under this paradigm, “computation” can be quite diverse — it can be storing and releasing energy periodically to stabilize locomotion (e.g., passive walker (?)), or conforming to complex objects to assist manipulation (e.g., vacuum jamming gripper (?)), or pre-processing sensory data to assist perception (e.g., bat ear that mechanically process the incoming sound waves to assist object localization (?)). Therefore, the physical computing defined in this paper can be an example of morphological computation, but it has a more structured definition. That is, many morphological computation examples will not be considered as physically computing in this paper because they do not have the “encoding-kernel evolution-decoding” architecture, and they are not reprogrammable.
On the other hand, a mechanical construct — e.g., architected materials or soft robotic body — that can incorporate encoding, decoding, and programming would meet the necessary condition to function as a physical computer. Moreover, based on these definitions, we will adopt the theory from Jaeger et al., and categorize physical computers into two sets (Figure 1b) (?, ?).
@@
One is analog, where input and output signals are continuous, and the evolution from input to output is governed by smooth (and frequently physics-based) functions. Albert Michelson’s harmonic analyzer (?) and our human neural system are classical examples of analog computers. In soft robots, this can be accomplished by exploiting their bodies’ nonlinear dynamic responses for physical reservoir computing. The other type of physical computer is the algorithmic, where the input and output take a discrete format, and the evolution from input to output is programmed via a set of abstract logical rules. Charles Babbage’s difference machine (?) and our omnipresent CMOS-based computing chips are classical examples of an algorithmic computer. In soft robots, this can be accomplished by, for example, an assembly of mechanical Boolean gates featuring elastic bistability (i.e., mechanical logic gates). Table 1 summaries and compares the different computing approaches from this perspective.
section computing approach working principle suitable tasks
2 analog oscillator use nonlinear mechanics to generate continuous and rhythmic signals (and motions) locomotion gait sequencing
3 analog reservoir extract nonlinear input-output mapping or complex information from high-dimensional body dynamics classification, multi-modal sensing, and locomotion control
4 algorithmic computer use mechanical logic gates to conduct discrete operations according to abstract logical rules user interfacing and control, behavior switching
Table 1: Summary of physical computing approaches discussed in this perspective.
@@ 2 Analog Oscillators and Rhythmic Motion
Figure 2: Analog oscillators and rhythmic deformation. (a) Electronics-free pneumatic control: A soft ring-oscillator circuit generates rhythmic leg actuation from a constant pressure input, enabling a quadruped to produce diagonal-couplet walking gaits. A bistable 4/2 switch selects gait direction, and dual oscillators set the phase between leg pairs (image adapted from  (?) with permission). (b) Controller-free SMA modular robot: A curved mono stable beam and a mechanical slider can convert a single DC power supply into sustained self-oscillation. And a bistable switch can alternate power supply between the front and back modules for out-of-phase deformation and crawling (adapted from (?), $C$ BY 4.0). (c) Twisted LCE ribbon robot: The ambient heating drives continuous self-rolling of this robotic structure for locomotion. When the robot contacts an obstacle, it will store elastic energy and then snap to reverse its direction, enabling autonomous avoidance and maze escape (adapted from (?), $C$ BY-NC-ND 4.0). Note that all scale bars are approximate.
An example of analog oscillator applied to soft robots: In the quadruped robot shown in Figure 2(a), rhythmic and coordinated leg swing motions are generated by an entirely pneumatic ring oscillator  (?). More specifically:
  • Input encoding: In this robot, a small pressure tank supplies a constant pressure ($P_{+}$) to drive the analog oscillator.
  • Kernel: The oscillator circuit is the kernel. The three soft valves inside this circuit serve as inverters with a built-in delay, and a snap-through membrane alternates between closed and open flow paths, allowing the high-pressure flow to advance around the ring. This mechanism essentially transforms the steady input into a phase-shifted sequence of pressure pulses at the three nodes.
  • Output decoding: The pulsed output pneumatic pressure from the oscillator flows to the corresponding soft legs, which convert the pressure inflation into mechanical swing motions for walking.
  • Re-programming: In addition, a soft bistable valve and tethered mechanical controller are added to swap two connections like a latching switch, so triggering the valve can reverse the output pressure pulse sequence, thus reversing the locomotion direction.
@@
Challenges and opportunities of analog oscillators: Analog oscillators are simple yet robust. They can tightly integrate with the soft robot’s body to generate locomotion without additional electronics. However, analog oscillators can suffer from programmability and scalability constraints: Their dynamics are hard-wired into physical design. That is, the oscillator geometry, mechanical architecture, and constitutive material properties fully determine the output frequency and phase pattern. As a result, “programming” the kernel’s input-output evolution might require re-design rather than a straightforward parameter tuning. In addition, as the number of oscillators increases, fabrication tolerances and material variability can introduce mismatches that degrade synchronization. One can address these limitations by integrating the oscillators with other, more easily programmable components (e.g., combining the oscillator with fluidic logic gates as we show later in Section 4), and using high-precision manufacturing techniques at the smaller physical scale (as we discuss later in Section 5).
@@ 3 Analog Physical Reservoir Computing
While an analog oscillator provides a promising alternative to micro-controllers for computing and generating rhythmic motions, its information processing capability is largely embedded in its physical architecture. Recent work has shown that architected mechanical and metamaterial-based systems can support multiple motion sequences through controlled switching of actuation frequency  (?, ?, ?), rather than through real-time algorithmic control. Nevertheless, the space of attainable behaviors in these systems remains discretely prescribed by design and reconfiguration pathways. In contrast, there is an emerging notion of informational embodiment, where the combinatorial richness of a soft robot’s body deformation encodes spatiotemporal patterns without relying on symbolic or centralized representations  (?). This concept bridges soft-body dynamics and computation, offering another pathway toward decentralized, analog computing.
@@
Figure 3: Open and closed-loop physical reservoir computing in soft robots. (a) Open-loop PRC for information perception: A modular manipulator with embedded strain gauges is driven by SMA actuators. Its high-dimensional body dynamics (measured by the strain $s_{i}(t)$) serve as reservoir states, which can be processed with trained linear readout $w_{i}$ to decode and identify the payload (image adapted from (?) with permission). (b) Other examples of information perception with open-loop PRC, including terrain classification(image adapted from (?) $C$ BY-ND 4.0), wind detection on a compliant membrane wing (image adapted from (?), $C$ BY 4.0), and a self-sensing shape memory alloy actuator that could predict its end effector trajectory (image adapted from (?) with permission). (c) Closed-loop PRC for embodied control: A quadruped robot uses its compliant spine as the reservoir. The four outputs of the reservoir kernel are fed back to the leg actuators to generate robust and adaptable locomotion gaits (image adapted from (?) with permission). (d) Other examples of control embodiments with closed-loop PRC, including manipulation with a multi-segment continuum arm (image adapted from (?) with permission) and a surface-swimming robot (adapted from (?), $C$ BY 4.0). Note that all scale bars are approximate.
@@
An example of open-loop reservoir applied to soft robots: The open-loop framework allows PRC to enhance the perception of soft robots by transforming their bodies into multi-modal computing sensors. A compelling demonstration of this is the modular manipulator equipped with SMA coil actuators and simple strain gauges (?) (Figure3a). The manipulator’s nonlinear body dynamics serve directly as the source for physical reservoir computing. More specifically:
  • Input encoding: When the manipulator grasps and lifts different payloads, its SMA actuators generate pulsed forces to “wobble” the body slightly.
  • Kernel: In this robot, the soft body itself is the kernel (or reservoir). As the SMA wobbles the manipulator and its payload, the resulting body vibration, denoted as $s_{i}(t)$, is captured by the strain gauges. Such a vibrational response is rich and nonlinear, so its spatiotemporal feature contains information about the weight and orientation of the payloads.
@@
  • Re-programming: The readout weights $w_{i}$ in the output layer are trained by regression methods, which can be adjusted according to the particular computing task at hand.
This example clearly illustrates how the open-loop reservoir computing framework enables soft robots to conduct spatiotemporal filtering through their intrinsic deformation dynamics, allowing them to extract complex information without requiring dense sensor arrays or extensive digital processing. As a result, numerous studies have emerged utilizing different soft robotic platforms. For instance, a fabric-based soft manipulator can estimate joint bending angle and payload weight simultaneously using only a few distributed pressure sensors  (?). Contact dynamics in a soft arm allow for tactile sensing and object property estimation without electronic skins  (?). For environment monitoring, a brush-like flexible sensor encodes surface textures through passive contact  (?)(Fig.3b, up), while in aerial applications, a flapping-wing robot detects wind direction directly from wing deformation, eliminating the need for airflow sensors  (?)(Fig.3b, middle). A SMA reservoir  (?) demonstrates the capability of predicting the future trajectory of its end effector under various driving signals (Fig. 3b, bottom). Collectively, these demonstrations show that soft robots with sparse, low-dimensional sensors can nonetheless achieve high-dimensional perception by exploiting their own body as the physical reservoir—making PRC a minimalist yet powerful strategy for embodied sensing.
@@
An example of closed-loop reservoir applied to soft robots: Beyond information perception, PRC enables soft robots to autonomously generate periodic and robust motions by embedding control into their intrinsic body dynamics. That is, instead of connecting robotic actuators to external digital controllers, one can feed the body reservoir’s output back to these actuators for real-time motor behavior control. In this case, the deformation dynamics of the robotic body and its interaction with the environment play a critical role. A representative example is the quadruped robot with its flexible spine serving as the reservoir computing kernel (?)(Fig. 3c). More specifically,
  • Kernel: In this robot, its flexible spine is the kernel (or reservoir). Its intrinsic body dynamics are rich and nonlinear, capable of projecting input signals into a high-dimensional state vector.
  • Input encoding and output decoding: In this case, the linear readout layer performs weighted linear summations, mapping the internal force and strain of the flexible spine into four control commands, each for a motored leg. The readouts are first trained in an open-loop setup with teacher forcing. Once the training is complete, the loop is closed. As a result, the reservoir computer’s output is also the input, eventually creating a self-sustained locomotion gait.
  • Re-programming: Once the loop is closed, and the robot can perform trotting, bounding, or turning — with a strong ability to recover from disturbance — simply by switching the readout weights. This embodied controller demonstrates how PRC converts compliant mechanics into an energy-efficient control system: the body remains unchanged, only the readout is “programmed,” and feedback routes it from perception to action.
@@
Challenges and opportunities of analog physical reservoirs: PRC offers an appealing framework to embody computation directly into soft robotic systems. In this framework, one can directly ”multipurpose” a soft robotic body into a reservoir without substantial redesign, and quickly switch the computing function by adjusting the readout weights. However, reservoir computing also faces several fundamental challenges. The first is repeatability. Real-world physical systems can not guarantee identical dynamic output across multiple experiments; slight variations in fabrication, boundary conditions, temperature, and material behavior, along with drift and aging, can shift the reservoir’s dynamic responses and degrade computing performance. A second challenge is noise, as real hardware inevitably introduces sensing and actuation noise that can be amplified by the reservoir’s nonlinear dynamics, leading to degraded or even unstable output. The third challenge involves scaling. One can always increase the number of reservoir states to improve performance, but it means more sensors, more wires, and heavier data-processing burdens. Possible improvements across these areas include better operating-point stabilization and calibration procedures, noise-aware training with improved sensing electronics, and more efficient sensing architectures or dimensionality-reduction strategies that capture the essential physical dynamics without overwhelming the hardware. These efforts collectively point toward more reliable, robust, and scalable physical reservoir computing systems for soft robotics.
4 Algorithmic Physical Computing and Mechano-Logic